Mark works strictly on commission of his gross sales from selling two different products for his company. Last​ month, his gross sales were ​$90,000. If he earns 5​% commission on product A and 3​% commission on product​ B, what were his gross sales for each product if he earned ​$4040 in total​ commission?
Answer by ikleyn(52776)

.
    x +     y = 90000    (1)

0.05x + 0.03y = 4040     (2)


Multiply equation (2) by 20 (both sides). Keep equation (2) as is. You will get


    x  +    y = 90000    (1')

    x  + 0.6y = 80800    (2')


Now subtract equation (2') from equation (1'). You will get

         0.4y = 90000 - 80800 

         0.4y = 9200

            y = 9200/0.4 = 2300.


ANSWER.  Gross sale for product B was  $23000;  gross sale for product A was the rest,  90000-23000 = 67000 dollars.


CHECK.  0.03*23000 + 0.05*67000 = 4040 dollars.    ! Correct !

Solved.

I used the Elimination method to solve the system of equations.